### v1

# Effective ML Model Versioning in Edge Networks

Abstract
Machine learning (ML) models, data and software need to be regularly updated whenever essential version updates are released and feasible for integration. This is a basic but most challenging requirement to satisfy in the edge, due to the various system constraints and the major impact that an update can have on robustness and stability. In this paper, we formulate for the first time the ML model versioning optimization problem, and propose effective solutions, including the automation with reinforcement learning (RL) based algorithm. Without loss of generality, we choose the edge network environment due to the known constraints in performance, response time, security, and reliability. The performance study shows that ML model version updates can be fully and effectively automated with reinforcement learning method as compared to other approaches. We show that with a carefully chosen range of traffic load values, the proper versioning can improve the security, reliability and ML model accuracy, while assuring a comparably lower response time.
I Introduction
The field of AI and ML has seen unprecedented advancements over the past decade, where AI models are continuously being improved in terms of size, performance, robustness and accuracy [1]. In this rapidly evolving domain, ML-based domain applications, such a industry and health care, need to be continuously updated. This comes in addition to related ML software updates as well as data updates. Recent ML Operations (MLOps) advancements recognize that automating and operationalizing ML products presents a new challenge [2]. To assure stability and manage applications and systems that use ML, an effective versioning of ML models, data and code is critical. Considering that most network management operations rely on ML based models, and most applications in IoT, cloud and edge computing today require ML-based functions, this challenge appears even more significant.
Model robustness is one of the key challenges in ML today, which requires creation of new versions of ML models [3]. The ML model updates on the other hand add performance overhead, and in general can decrease the stability of the system. In some domain applications, like health care, updating ML model is also a matter of safety [4]. In edge networks with constrained resources, several conflicting objectives need to be considered when updating ML models, such as constraints on accuracy, reliability, security, system stability, and resource utilization [5]. Adding any new and enhanced model is contingent upon modifying the existing system, and any changes fundamentally carry risks to destabilize the performance [6]. Only a few papers so far addressed this issue, like in [7] on how to update ML model for data series prediction, focused on a a single ML model. Today, multiple ML models updates need to be considered with various conflicting constraints.
In this paper, we focus on the problem of ML model versioning and updates in constrained networks. To this end, we formulate for the first time the ML versioning optimization problem and propose an RL-based decision making algorithm to automate and optimize the ML model update process. We also propose the related functional network management architecture of the ML model Update Manager in edge networks. We analyze the performance of our proposed approach by comparing it to conventional update methods used in today’s systems. The results show that under a certain network load, RL algorithm achieves a comparably faster response times, while improving model related parameters, such as accuracy, security and reliability. When the network load is higher, the automation with the RL algorithm becomes comparable to a random update process, guess process, which can be explained by the fact that under stressed system, the queue becomes constantly full and replicas remain alive during the whole runtime, not allowing the update agent for downgrading the deployed version to an older one.
The rest of the paper is organized as follows. Section II describes the reference architecture. Section III formulates the ML model versioning problem. Section IV describes the RL-based solution. Section V presents numerical results. Section VI concludes the paper.
Figure 1: a) Reference architecture, b) Versions representation of an ML model.
II Reference Architecture
Fig. 1.a) depicts the reference system architecture, which features a sample physical infrastructure. The sample network consists of four edge computing nodes (E1, E2, E3 and E4) connected by links (L1, L2, L3 and L4), along with the related worker machines (W1, W2, W3 and W4), similar to our previous work [8]. To control this infrastructure, we assume a container orchestrator, typically Kubernetes, an API gateway, a replication manager, and an Update Manager (our proposal). The system can run multiple IoT-based domain applications, as shown on the left hand side, such as vehicular, smart city, etc. The ML training is assumed in the cloud. As such, the entire system is an example of what is today known as IoT-edge-cloud continuum.
Applications communicate via API Gateway from the IoT sources via the use of HTTP requests. Let us assume that all application functions are ML applications, and that for each ML application, i.e., ML model, the replication manager decides the creation and removal of model replicas as well as the queuing of the received HTTP requests towards the updates. The nodes in the edge network are assumed to be edge devices that can host a number of dockerized containers, depending on the capacity. Each of these containers belongs to a cluster that is managed by the container orchestrator. As previously noted, the new versions of machine learning models are assumed to be trained in the cloud and subsequently received by the container orchestrator. This is out of scope of this paper, as we only focus on models.
We propose the architecture of Update Manager module, which in order to work, consists of a Telemetry tool, a Knowledge Base, and, in our proposal, an RL update agent. Telemetry is required to store the logs of all information coming from the container orchestrator, such as topology information and information about newly trained ML models. Knowledge Base is required to include a Version Repository database. Version Repository is expected to maintain a record of detailed information belonging to the various ML model versions, including data on the resources required, the reliability and security.
The RL update agent uses the information about newly available versions from the Version Repository database and the information regarding the current state of the topology from the Telemetry module to make update decisions. While Replication Manager takes scaling decisions by deciding to deploy a new replica of a certain ML model, Update Manager will decide whether to utilize the earliest, most stable version or the most recently developed version of an ML model. In the latter case, the RL update agent will examine the already deployed replicas. After executing a request, it will check the availability of a new version and determine whether the existing replica should be used to process the next requests or should be terminated to spawn a new replica with the newest version available. The RL update agent is trained to find the optimal policy for the decision process.
Fig. 1.b) illustrates the ML model versioning. A version is composed of main version $X$ and a subversion $Y$ to the version $X.Y$. The main version update alters the entire specification of a ML function, whereas a subversion update only incrementally enhances specific features, such as security update. We assume that a main version update for a particular ML model must be scheduled at a fixed time and is capable of modifying all ML model-specific values in a fixed, predefined manner. Subversion updates can, on the other hand, occur at a random time. In this paper, we consider three equally likely and practically relevant types of subversion updates: Security, Reliability and Accuracy updates.
III ML Model Versioning Problem Formulation
III-A Assumptions
The notation is provided in Table I. We consider a single master multi-worker deployment, including a set of edge nodes $\epsilon=\{E_{1},\ldots,E_{n},\ldots,E_{N}\}$, where $E_{n}$ represents the $n^{\text{th}}$ edge node. We assume that each edge node is constrained by a certain amount of capacity modeled as a number of CPU units $C_{n}$, RAM space $R_{n}$ and Disk space $D_{n}$. We consider a set of ML models denoted as $\kappa=\{1,\ldots,k,\ldots,K\}$. We assume that each ML model $k$, can have a main version $x\in[0,X]$, and a subversion $y\in[0,Y]$. We define a decision variable $a=1$ when updating a ML model $k$ of version $x.y$ to a newer version $[x+1].y$ or $x.[y+1]$, 0 otherwise. We define a mapping function $g(x,y,a)$ that gives the new version of an ML model $k$ of version $x.y$ after a decision update $a$. Each version $x.y$ of a ML model $k$ requires an amount of resources $b_{k,x,y}=[b_{k,x,y}^{\text{cpu}},b_{k,x,y}^{\text{ram}},b_{k,x,y}^{\text{disk}}]$, of CPU units, RAM and disk space. Each ML model $k$ of version $x.y$ has a average service rate $\mu_{k,x,y}$ and an average inter-arrival rate $\lambda_{k,x,y}$, assuming for simplification that $\mu_{k,x,y}=\mu,\lambda_{k,x,y}=\lambda,$ $\forall k,x,y$ We adopt an $M/M/N$ queuing system to analyze the system behavior and to compute the traffic load (in Erlang):
$\text{load}=\frac{\lambda}{N\cdot\mu}$ (1)
III-A1 ML Model Placement
Each ML model of class k requires an amount of resource that can be served by at least one available edge node, such that:
$b_{k,x,y}^{\text{cpu}}\leq\max_{\forall n\in[1,N]}\{C_{n}\},\quad\forall k,x,y\in[1,K]\times[0,X]\times[0,Y]$ (2)
$b_{k,x,y}^{\text{ram}}\leq\max_{\forall n\in[1,N]}\{R_{n}\},\quad\forall k,x,y\in[1,K]\times[0,X]\times[0,Y]$ (3)
$b_{k,x,y}^{\text{disk}}\leq\max_{\forall n\in[1,N]}\{D_{n}\},\quad\forall k,x,y\in[1,K]\times[0,X]\times[0,Y]$ (4)
We denote by $\delta_{k,x,y}(n)$ the number of replicas from ML model $k$ of version $x.y$ allocated in node $E_{n}$. The resource allocation has the capacity constraints, i.e.,
$\sum_{k=1}^{K}\sum_{x=1}^{X}\sum_{y=1}^{Y}b_{k,x,y}^{\text{cpu}}\delta_{k,x,y}(n)\leq C_{n},\quad\forall n\in[1,N]$ (5)
$\sum_{k=1}^{K}\sum_{x=1}^{X}\sum_{y=1}^{Y}b_{k,x,y}^{\text{ram}}\delta_{k,x,y}(n)\leq R_{n},\quad\forall n\in[1,N]$ (6)
$\sum_{k=1}^{K}\sum_{x=1}^{X}\sum_{y=1}^{Y}b_{k,x,y}^{\text{disk}}\delta_{k,x,y}(n)\leq D_{n},\quad\forall n\in[1,N]$ (7)
III-A2 Delays
We consider four types of delays: processing delay, transmission delay, delay to spawn a replica (where applicable) time and queuing delay. The processing delay $\tau^{p}_{k,x,y,n}$ of an ML model $k$ of version $x.y$ deployed through a pod in node $E_{n}$ is distributed around an average value. We consider transmission and propagation times to be a constant parameter since they depend on the distance between worker (edge) nodes and the so-called master node (corresponding to the practical case in Kubernetes). In a single master multi-worker deployment, the routing and path computation is typically managed by the container orchestrator (e.g., Kubernetes). We denote by $\tau^{t}_{k,n}$ the transmission delay of a request $f(k)$ between master node and worker node $E_{n}$. The spawn time $\tau^{s}_{k,n}(a)$ of a replica of ML model $k$ of version $x.y$ with a decision update $a$ is assumed to be constant for all type of ML models. Considering that the creation of a new replica is triggered by either the Replication Manager while scaling up the system, or by Update Manager, when we decide to update the deployed replica of the ML model $k$ with version $x.y$ to a newer version $x.[y+1]$ or $[x+1].y$. When the ML model $k$ is not updated, we set the spawn time $\tau^{s}_{k,n}(a=0)=0$.
We assume the system has a queuing buffer for each ML model $k$, assessed by Replication Manager, which can use a smart monitoring algorithm, such as the described in [9], to scale the system. We denote by $\tau^{q}_{f(k)}$ the total queuing delay of a request $f(k)$ of ML model $k$. The queuing delay is measured as the difference between the departure request time and the arrival request time plus the transmission plus the spawn time and the processing time. Finally, the total delay $\tau_{f(k)}$ for each request $f(k)$ of ML model $k$, processed using a replica that runs version $x.y$ is given as:
$\begin{split}\tau_{f(k)}(x,y,a)=&\tau^{p}_{k,g(x,y,a),n}+\tau^{t}_{k,n}+\tau^{s}_{k,n}(a)\\ &+\tau^{q}_{f(k)},\quad\forall k\in[1,K]\end{split}$ (8)
The average delay for all the requests, defined as an objective to minimize, is given by:
$\mathcal{O}_{1}(\Delta,a)=\frac{1}{KF}\sum_{k=1}^{K}\sum_{f=1}^{F}\tau_{f(k)}(x,y,a)$ (9)
where $\Delta$ describes the state of the network, including nodes, and deployed replicas of ML models and their versions.
III-A3 ML Model Accuracy
After training an ML model $k$, we obtain a new version that is assumed to have a higher accuracy. We denote by $Acc_{k}$($x,y,a$) the accuracy of a ML model $k$ with version $x.y$ after taking an update decision $a$. The average accuracy of all requests is given by:
$\mathcal{O}_{2}(\Delta,a)=\frac{1}{KF}\sum_{k=1}^{K}\sum_{f=1}^{F}Acc_{k}(x,y,a)$ (10)
III-A4 Security and Reliability
We assume that updates can improve the security and reliability of an ML model, defined as parameters $Sec_{k}(x,y,a),Rel_{k}(x,y,a)$ assigned to ML model $k$ with version $x.y$ after taking an update decision $a$. We define two objective functions for the security and reliability as an average among all requests given by:
$\begin{split}\mathcal{O}_{3}(\Delta,a)=\frac{1}{KF}\sum_{k=1}^{K}\sum_{f=1}^{F}Sec_{k}(x,y,a)\\ \mathcal{O}_{4}(\Delta,a)=\frac{1}{KF}\sum_{k=1}^{K}\sum_{f=1}^{F}Rel_{k}(x,y,a)\end{split}$ (11)
III-B Problem Formulation
We define our ML Model Versioning Problem (MMV) as a multi-objective optimization problem, aiming at maximizing the average ML model accuracy, security and reliability of all processed requests while minimizing the average delay. The MMV problem is formulated as follows:
[…]
Table I: Notations
Notation Description
$N$, $L$ Number of edge nodes and links
$C_{n}$, $R_{n}$, $D_{n}$ N. of CPU cores, RAM and disc in node $n$
$K$ Number of ML models.
$f(k)$ Request to ML model $k$
$b_{k,x,y}$ Required resources by ML model $k$
$X$ Number of main versions of an ML model
$Y$ Number of subversion of ML model
$\delta_{k,x,y}(n)$ Number of replicas of ML model $k$
with version $x.y$ in node $n$
$\tau^{f(k),k}$ Total delay of $f(k)$
$\tau^{p}_{k,g(x,y,a),n},\tau^{t}_{k,n},$ Processing, transmission, spawning, and queuing
$\tau^{s}_{k,n}(a),\tau^{q}_{f(k)}$ delay of $k$/$f(k)$ in node $n$, respectively
$\mu$, $\lambda$ Average service and arrival rates
$S$ Set of states representation for RL
$Z_{k}$ Set of resource availabilities for ML model $k$
$f_{k}$ Integer variable that defines the ML model $k$
$Q_{k}$ Queue size of queue for ML model $k$
$e$ Binary values to indicate the type of event
$\Omega$ Set of binary values indicating the type of update
$a(s)$ Set of RL actions
$R(s,a)$ Reward for RL
$w$ Weights of the RL Reward function
$\psi$, $\vartheta$, $\eta$, $\upsilon$ Delay, security, reliability, accuracy of the event
$\epsilon$ Exploration probability
$s^{\prime}$, $a^{\prime}$ , Next state and next action for RL
$\alpha$, $\gamma$ Learning rate and discount factor for RL
$Q_{t}(s,a)$ Q-Value of state $s$ and action $a$ for RL
IV Q-learning Based Update Decision Management
The previous problem formulation requires optimizations, which even for small networks, and multiple versions of ML models, and under multiple constraints, can be prohibitively complex. We therefore go immediately in media res of an RL based approach, and leave the optimization programming and the complexity analysis for future work. WIth RL, the system will make decisions when either: i) creating a replica during system scaling the system, or ii) processing of a request from the queue. The former concerns the deployment of a replica with either the most recent version of the ML model, or the initial version; the latter makes the decision on the basis whether the currently deployed replica should be utilized, or a newer version, if available, is to be deployed.
The RL model used here includes the commonly known key elements: agent, state space $S$, action space $A$, system reward $R$, and environment. RL agent computes the delays, security and reliability values associated with function requests, and assigns it a reward to the decision. RL model employs a Q-table to record, at each decision-making stage, the information required by the agent to make an informed choice. This includes the current state, the action to be taken, the associated reward, and the subsequent state. The system state $S$ at time $t$ for the RL model is represented by: resource availability, ML model id, queue length, event type and a list indicating the types of updates available for the ML model at stake:
$S=\{s\mid s=(Z_{k},f_{k},Q_{k},e,\Omega)\}$ (13)
where set $Z_{k}=\zeta_{1,k},...,\zeta_{n,k},...,\zeta_{N,k}$ encodes the nodes availability in a binary manner, and $\zeta_{n,k}$ is calculated from $\zeta_{n,k}=CPU_{n,k}\land RAM_{n,k}\land Disk_{n,k}$. If enough resources are available for a deployment of ML model $k$ in node $n$ $CPU_{n,k}$ will be 1; otherwise 0. The same principle applies for $RAM_{n,k}$ and $Disk_{n,k}$. Consequently, the value of $\zeta_{n,k}$ will be 1 if there are enough resources available for CPU, RAM and Disk to deploy the ML model. Therefore, the value of $\zeta_{k}$ indicates whether ML model $k$ can be deployed on node $n$. The ML model identification $f_{k}$ is an integer variable that defines the type $k$ of a ML model. $Q_{k}$ is an integer variable that defines the queue size for this ML model. A binary variable $e$ is used to indicate the type of event that occurs in the system. In the case of an arrival event, the value of $e$ is set to zero, while in the case of a departure event, the value of $e$ is set to one. The set, noted as $\Omega$ is a list of length four, in which each element is a binary value $\Omega=[\omega_{1},\omega_{2},\omega_{3},\omega_{4}]$. Each $\omega$ represents a distinct type of update that differentiates the two ML model versions under consideration: main version, subversion security, subversion reliability, or subversion accuracy updates.
The RL Update Manager has a binary set of actions $a(s)$ to take, whenever a new version and one of the following events: a) the replication decision algorithm decides to create a new replica, or b) replica $m$ is processed from a queue. The action variable is defined as:
$a(s)=\begin{cases}1&\text{if instance }m\text{ is updated}\\ 1&\text{if new instance }n\text{ is created with newest version}\\ 0&\text{otherwise}\end{cases}$ (14)
At each system state and following the execution of an action $a(s)$, the update decision agent receives a reward. The reward function $R(s,a)$ used in our update decision agent considers the total delay, in addition to the values of security, reliability and accuracy. The agent will select actions that minimize the delay while simultaneously optimizing security, reliability and accuracy. The reward is defined as follows:
$R(s,a)=-w_{1}\psi+w_{2}\vartheta+w_{3}\eta+w_{4}\upsilon$ (15)
where $s$ is a state, $a$ is an action, $\psi$ is the total delay for the processed event, $\vartheta$ is the security value, $\eta$ is the reliability values and $\upsilon$ is the accuracy values of this event.
Algorithm 1 RL-based Update Decision Algorithm
1:Input: events, $b_{k}$, $\mu_{k}$, network state ($Z$, $Q$)
2:Initialization: $\mathcal{Q}_{t}$, $\epsilon$, $\alpha$, $\gamma$, episodes, decay
3:for each event do
4:    if update available then
5:        $s$leftarrow ($Z$, $Q$, $e$, $f$),     $a$ = UpdateDecision($s$)
6:        $Z^{\prime}$, $Q^{\prime}$, $e^{\prime}$, $f^{\prime}$leftarrow Find next state parameters
7:        $s^{\prime}$leftarrow ($\Delta^{\prime}$, $Q^{\prime}$, $e^{\prime}$, $f^{\prime}$)
8:        $\mathcal{Q}_{t}$leftarrow UpdateQTable($s$, $a$, $s^{\prime}$)
9:        $s$leftarrow $s^{\prime}$
10:    else
11:        pass
12:    end if
13:    if $\epsilon$ $>$ $\epsilon_{\text{min}}$ then
14:        $\epsilon$leftarrow decay $\cdot\epsilon$
15:    end if
16:end for
The goal of the Algorithm 1 is to select an action, as defined in Eq. 14. During warm-up, where previous actions were taken, or the RL model has not learned yet to accurately choose actions, an exploration phase is needed. We consider an $\epsilon$-greedy approach to explore the search space and try new actions. We denote by $\epsilon$ the exploration probability, where at each step the RL agent chooses randomly an action with probability $\epsilon$ and uses the accumulated knowledge with probability $1-\epsilon$. As the simulation progresses, the exploration-exploitation factor declines linearly until it reaches a minimal value. At each point of decision, the agent checks the state of the environment from Eq. 13 to take an action a using $UpdateDecision(s)$ function, which chooses an available action with the highest q-value, considering the $\epsilon$-greedy approach.
V Numerical Evaluation
ML Model id 1 2 3 4 5
Version 0 to 10
Subversion 0 to $\sim$ 10000
Avg service time 10
Avg inter-arrival time 3.25 - 8
CPU req 1 2 3 4 5
RAM req 1 1 2 2 6
Disk req 0.01
Spawn time 10
Reliability 0.9 - 1
Security 0.6 - 1 0.6 - 1 0.6 - 1 0.7 - 1 0.6 - 1
Accuracy 0.5 - 1 0.5 - 1 0.5 - 1 0.7 - 1 0.7 - 1
Table II: ML model parameters
Node
CPU Capacity
(in CPU cores)
RAM Capacity
(in GB)
Disk Capacity
(in TB)
Transmission time
to master
1 16 16 1 0
2 16 16 1 2.75
3 16 16 1 7.25
4 16 16 1 10.25
Table III: Deployed topology parameters
The objective of our numerical evaluation is to assess the performance of the RL update agent in determining whether to proceed with an update to the newest version of an ML model, or to maintain operation of the existing stable ML version. To this end, we used an event based simulator running up to 2 million arrival request events and around 5000 version update events, which assured a 98% confidence. We used a First-fit algorithm for resource allocation, a monitoring-based scaling for creating and removing replicas, which simply replicate whenever a the load reach a certain threshold, and remove replicas when the queue gets empty, the allocation and scaling were well analyzed in [9], and theoretically in [10]. We compare different types of update strategies in order to evaluate the performance of the RL update agent: i) always update, ii) never update and iii) select an update randomly. Our topology consists of 4 nodes in a star topology around the master node of a cluster, which in our case is node 1, see Table III, considered as a worker node as well. All nodes are assumed to have 16 CPU cores, 16 GB of RAM, and 1 TB of disk capacity. The transmission delay between each node to the master node is assumed to be 0, 2.75, 7.25, and 10.25 (ms).
We consider 5 distinct ML models to be deployed in the system, as outlined in Table II. The main versions of each ML model vary from 1 to 10, subversions from 0 to 104, with average service time $\mu=10$, average inter-arrival time from 3.25 to 8, the spawn time is 10 ms, for all ML models. We assume that the CPU requirement is different between models, $k$ CPU units for ML model k. The RAM requirements is 1 BG for model 1 and 2, and 2 GB for model 3 and 4, and 6 for model 5. The disk req. is the same for all models with 0.1 TB. We set a security parameter per version between 0.6 and 1 for model 1, 2, 3, and 5, and between 0.7 and 1 for model 5. The reliability parameter varies between 0.9 and 1 for all models, and the model accuracy between 0.5 and 1 for model 1, 2, and 3, and between 0.7 and 1 for model 4 and 5. A subversion update will randomly alter the values of reliability, security or accuracy by a small margin. For the reward function defined in eq. (15), the weights were chosen to be $w_{1}=1$ , $w_{2}=10$ , $w_{3}=10$ and $w_{4}=10$. For Q-table update, we set $\alpha=0.01$, $\gamma=0.99$, the initial exploration probability $\epsilon=1$ , this will decay with an $decay$ up to an minimal $\epsilon_{\text{min}}=0.001$ which will be reached after half of the scheduled events have been processed. The RL settings are mostly used for training and proved a good performance.
(a) Average Security
(b) Average Reliability
(c) Average Accuracy
(d) Average Delay
Figure 2: Performance metrics with various loads of different update approaches.
(a) Always Update
(b) Never Update
(c) Random Update Decision
(d) RL Update Decision
Figure 3: Used subversion over events of different update approaches for Load of 31.5% in an interval for ML model 5
Fig. 2 illustrates the performance metrics for various update approaches under different traffic load conditions using Eq. 1, varying from 30% up to a maximum value of about 80%. Figs. 2(a), 2(b), 2(c) display the average security ($\mathcal{O}_{3}$), average reliability ($\mathcal{O}_{4}$), and average accuracy ($\mathcal{O}_{2}$) as a function of the system load. The performance achieved follow follow the same trend. As the traffic load increases, the parameters for all update policies are observed to decrease. As anticipated, the method always update is performing the best, as it consistently selects the most recent version with the best parameters (security, reliability, accuracy), i.e., with latest security updates. Nevertheless, it decreases with increasing traffic load, because the higher the traffic load the less updates received and processed by the system for the same number of requests. For low traffic load, RL is the second-best approach after always update, as it learns to enhance the parameter values. For higher traffic loads, RL behaves similar to random update, because when the traffic load is high, the system never scales-down replicas, which prevents the system from using old and stable versions and there is minor variation in the decision space. The performance of never update remains consistent for all system loads, as the system maintains identical versions of all ML models.
Figure 2(d) shows the average delay for all update strategies. As expected, the delay increases with increasing load. Never update performs the best, which avoids the additional spawn time periods. In order to update a deployed ML model, we shut-down the existing replica and spawn a new one with the most recent ML model, which adds additional delays. RL achieves a lower delay than the random for lower traffic loads. For higher loads, it again exhibits behavior similar to random. As expected, always update performs comparably worse, due to the additional delays at each possible update.
We also illustrate of the evolution of the deployed subversion for the various decision models, see Figure 3. We show an interval of 1,000 events (after a phase of RL exploitation, and for that we plot the versions after 2 106 events), showcasing the deployed subversion for ML model 5. Always update (Figure 3(a)) shows the progressive increase in subversion over time. The minor regressions are due to a not-yet updated replicas for the same ML model, while other replicas have finished the update faster. Never update (Figure 3(b)) always deploys the oldest version and does not undergo any updates. Random update (Figure 3(c)) shows the decision of either to update a replica or retain the existing version, whereas when a new replica is created, the option of selecting a replica with subversion 0 is made randomly with no clear trend, which explain the behavior of the curve. Finally, RL strategy (Figure 3(d)) illustrates a combination of always update and randomly select an update trajectories, with proportions that maximizes the defined reward.
VI Conclusion
In this paper, we addressed the problem of automating ML model versioning in constrained edge networks, considering different kind of updates to ML models. We compared a RL based approach to other update strategies, which today are common practice. The simulation results demonstrate that the quality of RL-based decision depends on the traffic load. For less load, RL achieves comparable results to the always update strategy. At the same time, it outperforms the random update in all measurements. Perhaps most importantly, our study showed important limitations and thus directions for further research. We assumed that subversion updates always improve (linearly or exponentially), which is practically not always the case. Also we did not take the stability as a values for the optimization problem into consideration yet, which also is premium in real world systems. Also many more, and larger topologies need to be studied, along with the proper optimizations of the solution. Finally, we did not take into account any downgrading of a version. These and many other interesting findings are subject of future research.
Acknowledgment
This work was partially supported by the EU project MANOLO under GA no. 101135782.